# NOBLE - Neural Operator with Biologically-informed Latent Embeddings to Capture Experimental Variability in Biological Neuron Models

### v2

@@ 5 Conclusion
We introduced NOBLE for learning the nonlinear somatic dynamics across a population of HoF models for a single neuron. Rather than training separate surrogates for each bio-realistic model, NOBLE learns a single neural operator that captures the inherent variability observed in experimental neuron recordings by mapping biologically interpretable embeddings to voltage responses from current injections. Demonstrated on HoF models of a PVALB neuron, NOBLE correctly captured the diverse neuron dynamics with a 4200× speedup over traditional solvers, while maintaining accuracy across key electrophysiological features. NOBLE also allows for generating novel, bio-realistic neuron models through interpolation in the latent space, which was infeasible directly with HoF models. This allows for accurate predictions beyond the original set of HoF models, enabling faster model exploration and ensemble prediction. NOBLE’s interpretable latent space also offers new insights on how neuron characteristics affect neuron dynamics. While NOBLE successfully captures the dynamics across HoF models, it is currently limited to modeling single neurons. A natural extension of this work is to generalize NOBLE to multi-neuron settings, enabling applications such as neuron classification and predicting multi-neuron dynamics. Also, the embedding space used in this our experiments is low-dimensional, constructed from two interpretable features derived from a model’s F-I curve. Although NOBLE demonstrates strong performance in this configuration, extending it to a learnable, higher-dimensional continuous embedding space remains an exciting direction for future work.
@@ Acknowledgements
L.G. was responsible for the complete technical implementation of this work. C.A.A. and A.A. conceptualized this work. L.G., V.D., and B.T. jointly developed the methodology of the novel NOBLE framework. P.H.W. and C.A.A. provided neuroscience-specific domain expertise, contextualizing the relevance and impact of this work within the broader field of neuroscience. P.H.W. supplied the biophysical PDE models and neuron morphologies. L.G. and P.H.W. produced the figures. L.G., V.D., B.T., P.H.W., and C.A.A. co-wrote the manuscript. C.A.A. and A.A. provided supervision and insightful editorial comments.
@@
A.A. is supported by the Bren Endowed Chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 Senior Fellow program at Schmidt Sciences. C.A.A. is being supported by the National Institutes of Health R01 - NS120300 and R01 - NS130126. P.H.W. is being supported by the National Institutes of Health R01 - NS130126.
Appendix A The Fourier Neural Operator Architecture
Neural operators compose linear integral operators $\mathcal{K}$ with pointwise non-linear activation functions $\sigma$ to approximate highly non-linear operators. More precisely, we define the neural operator
where $\mathcal{P}$, $\mathcal{Q}$ are the pointwise neural networks that encode the lower dimension function into a higher dimensional space and vice versa. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters $\theta$ consists of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. [46] maintain a comprehensive open-source PyTorch library for learning neural operators, which serves as the foundation for our implementation. A Fourier neural operator (FNO) [22] is a neural operator using Fourier integral operator layers, which are defined via
where $R_{\phi}$ is the Fourier transform of a periodic function $\kappa$ parameterized by $\phi$. On a uniform mesh, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform (FFT). Here is a depiction of the Fourier Neural Operator (FNO) [22]:
Appendix B Impact of Subsampling on Neuron Features
We present the results of the analysis conducted to determine the maximum subsampling factor that preserves the fidelity of extracted neuron features, mentioned in Section 3.1. The results are displayed in Figures 9 and 10 for the low-pass filtering followed by decimation in time subsampling strategy.
This study shows that for most features, subsampling introduces negligible additional error. The most sensitive features were AP1_Width and AP1_Peak, which exhibited noticeable deviations at higher subsampling rates. To ensure we remain within the bounds of the intrinsic simulation-experiment discrepancy, we adopt a conservative downsampling factor of 3×3$\times$3 ×, which maintains fidelity while reducing computational load.
Appendix C Input Current Amplitude Distribution
Figure 11: Distribution of square-pulse amplitudes in $[-0.11,0.28]\mathrm{nA}$ considered. There is a spiking threshold (between 00 to $0.05\mathrm{nA}$) where neuron responses transition from being non-spiking to spiking.
Appendix D Neighborhood Considered in Interpolation Experiments
Figure 12: Latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of $\{\text{HoF}_{i}^{\textit{train}}\}$ (black dots) and $\{\text{HoF}_{i}^{\textit{test}}\}$ (red crosses) models. The latter lie in the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}_{i}^{\textit{train}}\}$ models. In the interpolation experiment of Section 4.3, we sample 50 models in a small neighborhood around a given model $\text{HoF}_{k}^{\textit{test}}$. We plot the boundary of this neighborhood (the blue ellipse) for an example of model $\text{HoF}_{k}^{\textit{test}}$. Note that this neighborhood does not contain any model used for training, and in particular none of the 50 new models sampled in that neighborhood have been used for training.
Appendix E Electrophysiological Features
@@
We list below 15 important electrophysiological features and metrics of interest when constructing neuron models (where AP denotes action potential and AHP denotes after-hyperpolarization):

### v3

@@ 5 Conclusion
We introduced NOBLE, a neural operator framework for learning the nonlinear somatic dynamics across a population of HoF models for a single neuron. Rather than training separate surrogates for each bio-realistic model, NOBLE learns a single neural operator that captures the inherent variability observed in experimental neuron recordings by mapping biologically interpretable embeddings to voltage responses from current injections. Demonstrated on PVALB and VIP neurons, NOBLE correctly captured the diverse neuron dynamics with a 4200× speedup over traditional solvers, while maintaining accuracy across key electrophysiological features. Importantly, our work is among the first to benchmark a deep learning based method for predicting membrane potential responses to intracellular current injections against experimental data from the human cortex. NOBLE also allows for generating novel, bio-realistic neuron models through interpolation in the latent space, which is not feasible with HoF models. This allows for realistic and efficient ensemble predictions beyond the original set of HoF models. NOBLE’s interpretable latent space also offers new insights into how neuron characteristics affect neuron dynamics. NOBLE opens a pathway toward modeling larger-scale brain circuits and leveraging multimodal latent spaces to determine relationships between gene expression, electrophysiology, and morphology, as discussed in Appendix˜F.
@@ Acknowledgements
L.G. was responsible for the complete technical implementation of this work. C.A.A. and A.A. conceptualized this work. L.G., V.D., and B.T. jointly developed the methodology of the novel NOBLE framework. P.H.W. and C.A.A. provided neuroscience-specific domain expertise, contextualizing the relevance and impact of this work within the broader field of neuroscience. P.H.W. supplied the biophysical PDE models and developed the multi-objective optimization pipeline. L.G. and P.H.W. produced the figures. L.G., V.D., B.T., P.H.W., and C.A.A. co-wrote the manuscript. C.A.A. and A.A. provided supervision and editorial comments.
@@
A.A. is supported by the Bren Endowed Chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 Senior Fellow program at Schmidt Sciences. C.A.A. is supported by the National Institutes of Health R01 - NS120300 and R01 - NS130126. P.H.W. is supported by the National Institutes of Health R01 - NS130126.
Appendix A Related Works in Machine Learning for Single-Cell Electrophysiology
Earlier applications of machine learning to single-cell electrophysiology focused on directly learning the dynamics of canonical point-neuron models like FitzHugh-Nagumo [30, 31] and Hodgkin-Huxley [32]. Fully connected and convolutional neural networks were trained to reproduce FitzHugh-Nagumo dynamics [33], while ResNet-based multilayer perceptrons showed promise in learning Hodgkin-Huxley dynamics [34, 35]. Physics‑informed neural networks (PINNs) extend conventional neural networks by introducing prior knowledge about the underlying dynamical system [36, 37, 38]. This formulation was used to predict FitzHugh-Nagumo dynamics [39] and to learn the Hodgkin-Huxley model ionic conductances from simulated voltage recordings [40]. Further refinements with PINNs incorporated wavelet bases to capture localized multiscale dynamics and compute derivatives analytically, improving both accuracy and convergence speed when training on the FitzHugh-Nagumo model [41]. While these methods demonstrate capabilities in capturing the core spiking dynamics of these simplified models, their ability to accurately represent the full spectrum of electrophysiological behavior, particularly the highly nonlinear onset of firing, remains largely untested. Moreover, as function approximators, they necessitate retraining for each new input stimulus, significantly limiting their practical utility.
To address some of these limitations, Centofanti et al. [42] explored using operator learning approaches for forward simulations of the Hodgkin-Huxley model. Among other approaches, FNOs showed promising results by demonstrating a strong capacity for learning the governing operator of this biophysical system. However, this work still exhibits key limitations and a limited scope: (1) it relies on relatively simple simulated data from a point-neuron model, (2) it does not explicitly attempt to capture the full spectrum of electrophysiological dynamics, particularly the highly nonlinear onset of firing, and (3) its formulation on a single operator inherently lacks the capacity to represent the trial-to-trial variability observed in biological recordings. Related work such as NeuPRINT [43] also leverages deep learning to capture trial-to-trial variability. However, it operates on slower in vivo 2-photon calcium imaging data and models population-level fluorescence dynamics, rather than the fast intracellular voltage dynamics of individual neurons.
Appendix B Electrophysiological Features
@@
We list below 16 important electrophysiological features and metrics of interest when constructing neuron models (where AP denotes action potential and AHP denotes after-hyperpolarization):
@@
  • mean_frequency: The mean frequency of the firing rate
@@
Appendix C Method
C.1 Impact of Subsampling on Neuron Features
We present the results of the analysis conducted to determine the maximum subsampling factor that preserves the fidelity of extracted neuron features, mentioned in Section˜3.1. The results are displayed in Figures 6 and 7 for the low-pass filtering followed by decimation in time subsampling strategy.
This study shows that for most electrophysiological features, subsampling introduces negligible additional error. The most sensitive features were AP1_Width and AP1_Peak, which exhibited noticeable deviations at higher subsampling rates. To ensure we remain within the bounds of the intrinsic simulation-experiment discrepancy, we adopt a conservative downsampling factor of $3\times$, which maintains fidelity while reducing computational load.
C.2 Input Current Amplitude Distribution
Figure 8: Distribution of square-pulse amplitudes in $0.11,0.28]\mathrm{nA}$ considered. There is a spiking threshold (between 0 to $0.05\mathrm{nA}$) where neuron responses transition from being non-spiking to spiking.
C.3 The Fourier Neural Operator Architecture
Neural operators [25, 26, 27] are a principled way to generalize neural networks to learn operators mapping functions to functions, with a universal operator approximation property [28]. Neural operators compose linear integral operators $\mathcal{K}$ with pointwise nonlinear activation functions $\sigma$ to approximate highly nonlinear operators. More precisely, we define the neural operator
where $\mathcal{P}$, $\mathcal{Q}$ are the pointwise neural networks that encode the lower dimension function into a higher-dimensional space and vice versa. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters $\theta$ consists of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. [55] maintain a comprehensive open-source PyTorch library for learning neural operators, which serves as the foundation for our implementation. Prior knowledge of the relevant physics laws and differential equations can also be incorporated as additional loss terms during training, to supplement or replace reference data, as done with physics-informed neural operators [57, 58, 59].
A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [29], and successfully applied to a wide range of problems [60, 61, 62]. A FNO is a neural operator using Fourier integral operator layers, which are defined via
where $R_{\phi}$ is the Fourier transform of a periodic function $\kappa$ parameterized by $\phi$. On a uniform mesh, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform (FFT).
Appendix D Experiments
D.1 Dataset
Figure 10: Latent representations of neuron models in the normalized ($I_{thr},s_{thr}$)-space. Black dots indicate the 50 training HoF models $\{\text{HoF}^{\textit{train}}\}$, and red crosses the 10 test HoF models $\{\text{HoF}^{\textit{test}}\}$ excluded during training.
We have access to 60 HoF models $\{\text{HoF}\}$ obtained using a multi-objective evolutionary optimization strategy. We use 50 of them during training, $\{\text{HoF}^{\textit{train}}\}$, and keep the remaining 10 $\{\text{HoF}^{\textit{test}}\}$ for testing.
Figure˜10 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of these HoF models.
The training dataset is composed of 75,600 samples, where the current injections are sampled as described in Section˜3.2, each of which is associated randomly to one of $\{\text{HoF}^{\textit{train}}\}$. The samples are generated using a numerical solver [44] built on the NEURON simulation environment [45].
D.2 Evaluation and Evaluation Metrics
NOBLE is trained using the relative L4 error, computed via
$\text{Relative Lp error}(x,y\,;\epsilon)=\frac{\|x-y\|_{p}}{\|y\|_{p}+\epsilon}=\frac{(\sum_{i}|x_{i}-y_{i}|^{p})^{1/p}}{(\sum_{i}|y_{i}|^{p})^{1/p}+\epsilon}.$ (4)
The relative L4 error was selected as the training loss after a preliminary training study on a small dataset, where it consistently preserved spike-related features, especially amplitudes and widths, more effectively than the relative L2 error. While this choice ensured the model captured the electrophysiological details most relevant to our setting, NOBLE is compatible with any other loss function that may be more suitable in different contexts.
Although NOBLE is trained to minimize the relative L4 error, we report results using the relative L2 error, as it provides a more common and interpretable measure of accuracy. We first evaluate performance on voltage traces. Note that even small temporal shifts between predicted and ground-truth voltage responses can result in large relative errors. To better evaluate physiologically meaningful behavior, we also report errors on four key electrophysiological features:
  • spikecount: number of spikes in the trace
  • AP1_width: width of the first spike at half amplitude
  • mean_AP_amplitude: mean amplitude of all action potentials
  • steady_state_voltage: average voltage after the stimulus
For benchmarking, we compare NOBLE predictions against numerical simulations obtained from HoF models and experimental data. HoF models were optimized to produce the closest approximations to experimental recordings and capture the biological variability required for a meaningful benchmark. In contrast, existing machine learning methods, such as the ones discussed in the introduction, are not designed to reproduce such variability and are therefore unsuitable as baselines. Furthermore, since the PDE solvers used to generate the dataset are themselves approximations with non-negligible error, driving prediction error below the solver–experiment gap risks overfitting to the solver rather than improving alignment with real recordings. Therefore, we tuned hyperparameters only up to the solver–experiment error level, as marginal gains on PDE data are unlikely to yield meaningful improvements relative to experimental recordings.
D.3 Testing on VIP Neuron HoF Models Included in the Training Set
To assess NOBLE ’s ability to perform well on different neuron models, we trained it on a VIP neuron using the same architecture as in the PVALB case with 1.8M trainable parameters. The model was optimized for 450 epochs with the Adam optimizer with learning rate 0.004 and the ReduceLROnPlateau scheduler with factor 0.8 and patience 4, minimizing the relative L4 error. For the embeddings, the neural operator in NOBLE takes the stacked embeddings of the normalized model features $I_{\textit{thr}}$ and $s_{\textit{thr}}$ associated with $\text{HoF}_{\ell}$ (with $K=1$ frequency) and $I$ (with $K=11$ different frequencies).
The trained model achieves a relative L2 error of 2.5% on voltage traces and relative L2 errors of 9.0% for spikecount, 20% for AP1_width, 10% for mean_AP_amplitude, and 0.99% for steady_state_voltage.
These results are comparable to those obtained for PVALB, indicating that NOBLE, with the same latent embedding space, also performs well when trained on different neuron types.
D.4 Neighborhood Considered in Interpolation Experiments
Figure 11: Latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of $\{\text{HoF}^{\textit{train}}\}$ (black dots) and $\{\text{HoF}^{\textit{test}}\}$ (red crosses) models. The latter lie in the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}^{\textit{train}}\}$ models. In the interpolation experiment of Section˜4.3, we sample 50 models in a small neighborhood around a given model $\text{HoF}_{k}^{\textit{test}}$. We plot the boundary of this neighborhood (the blue circle) for an example of model $\text{HoF}_{k}^{\textit{test}}$. Note that this neighborhood does not contain any model used for training, and in particular, none of the 50 new models sampled in that neighborhood have been used for training.
D.5 Auxiliary Ensemble Prediction Figures
Figure 12: Distribution of somatic voltage traces across HoF models $\{\text{HoF}^{\textit{test}}\}$ for current injections of $0.1\mathrm{nA}$ (top row) and $0.11\mathrm{nA}$ (bottom row). A) Ground truth voltage responses from HoF simulations, B) NOBLE predictions.
Figure 13: Distribution of somatic voltage traces across $50$ synthetic HoF models sampled from a small circle centered on a HoF model in $\{\text{HoF}^{\textit{test}}\}$ not experienced during training, as shown in Figure˜11. Results are shown for current injections of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom). The ground truth voltage response from the HoF simulation not experienced during training is shown in blue, and the $50$ synthetic NOBLE predictions in orange.
D.6 On the HoF Models Availability Requirement
Training NOBLE on 75,600 samples required approximately 4 days on a 64GB NVIDIA Tesla P100 GPU (300 epochs). Once trained, for a given input current, NOBLE can synthesize arbitrarily many voltage responses of synthetic HoF models almost instantaneously by interpolating within $\mathcal{CH}_{\text{{train}}}$ (Figure˜10). Thus, NOBLE amortizes the high upfront cost of bio-realistic neuron model generation and enables scalable response synthesis at negligible inference cost. To investigate how much model diversity is needed during training for effective amortization, we varied the number of HoF models used to construct the training set $\{\text{HoF}^{\textit{train}}\}$ while keeping the total number of samples fixed. Results are summarized in Table˜2 in terms of the relative L2 error of predicted voltage traces and the four key electrophysiological features when evaluated on the $\{\text{HoF}^{\textit{test}}\}$ models.
Table 2: Predictive performance of NOBLE on voltage traces and the four key electrophysiological features using the relative L2 error metric when the training set $\{\text{HoF}^{\textit{train}}\}$ is constructed by including varying numbers of models. Results are evaluated on $\{\text{HoF}^{\textit{test}}\}$.
#HoFs included Voltage Steady state voltage Spikecount AP1 width Mean AP amplitude
50 11.7% 2.0% 9.2% 350% 14%
40 10.9% 1.8% 19% 920% 17%
30 10.9% 1.9% 20% 3004% 20%
20 10.6% 1.9% 45% 1698% 20%
The relative L2 error on voltage traces is largely insensitive to HoF diversity, whereas spike-related features, particularly spikecount and AP1_width, degrade substantially as diversity decreases. As discussed earlier, AP1_width is particularly sensitive to small misalignments in spike peak and after-hyperpolarization minima, which can shift the half-voltage reference and lead to large relative errors even when the underlying traces are close. This makes AP1_width less reliable for direct comparison than other features. Overall, these findings indicate that while NOBLE amortizes the cost of HoF generation, effective bio-realistic synthesis still requires sufficient model diversity to learn robust representations of neural dynamics. Depending on the purpose of the study and the solver–experiment error gap, the number of HoF models required for training can be adjusted accordingly. Moreover, for any given electrophysiological feature of particular interest, further fine-tuning can be used to refine predictions and improve generalization, as we discuss next.
D.7 Additional Information on Fine-Tuning
Suppose the objective is to capture overall neural dynamics while placing particular emphasis on one specific electrophysiological feature. In this setting, the loss function can be designed to prioritize the feature of interest. Let $F$ denote a feature computed on the ground-truth signal and $\hat{F}$ the corresponding feature computed from NOBLE ’s output. A feature-specific loss can then be defined as $\mathcal{L}_{F}=\|F-\hat{F}\|$, which directly penalizes deviations in the feature of interest.
To illustrate this, we fine-tune the previously trained NOBLE to enhance sag_amplitude predictive performance. This feature is particularly relevant as it reflects the presence of the hyperpolarization-activated cation channel (Ih). In human cortex, the expression of Ih varies with cortical depth, making sag_amplitude a useful marker. Although the broader role of Ih in shaping neuronal and network properties is not yet fully understood, it is thought to regulate neural excitability and coincidence detection.
We start from a pretrained NOBLE model and further optimize the weights of the neural operator using the feature-specific loss. Relying solely on a feature-specific loss, such as $\mathcal{L}_{\text{sag}}$, risks causing NOBLE to overfit to sag_amplitude, improving that metric while degrading performance on other features and overall voltage trace fidelity. To mitigate this issue, one possible strategy is to introduce an anchor loss, as proposed in PINO [57], which penalizes deviations from the pretrained operator during fine-tuning. Combining the anchor loss with $\mathcal{L}_{\text{sag}}$ could constrain optimization so that gains on a single feature do not come at the expense of overall signal fidelity, thereby encouraging balanced gains across all electrophysiological features.
Another approach, which we adopt, is to define a composite loss that encourages further accuracy on the voltage traces while prioritizing the feature of interest, sag_amplitude:
$\mathcal{L}(\lambda)=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{sag}}.$
To evaluate this approach, we construct a smaller dataset of 19,600 stimulus waveforms with negative non-zero amplitudes, since this regime elicits non-firing responses for which the sag_amplitude can be reliably computed. We then fine-tune the pretrained NOBLE model by minimizing the relative L4 error for $\mathcal{L}_{\text{data}}$ and the relative L2 error for $\mathcal{L}_{\text{sag}}$, using the Adam optimizer with learning rate of $0.0005$, and the ReduceLROnPlateau scheduler with factor $0.4$ and patience $12$. The results, reported in Table˜3, summarize the relative L2 error of predicted voltage traces and sag_amplitude on the test set of $\{\text{HoF}^{\textit{train}}\}$.
Table 3: Predictive performance of NOBLE fine-tuned on sag_amplitude. Metrics are reported as relative L2 errors on voltage traces and on sag_amplitude, with the training set $\{\text{HoF}^{\textit{train}}\}$ constrained to samples with negative non-zero amplitudes. Results are evaluated on the test set of $\{\text{HoF}^{\textit{train}}\}$. Here, $\lambda$ denotes the weighting factor of the feature-specific loss in the composite loss.
Before optimization Epoch $100$ Epoch $200$ Epoch $300$
$\lambda$ Voltage Sag amplitude Voltage Sag amplitude Voltage Sag amplitude Voltage Sag amplitude
0 0.14% 69.6% 0.064% 22.2% 0.041% 20.3% 0.043% 19.2%
25 0.14% 69.6% 0.055% 13.2% 0.047% 12.2% 0.035% 9.6%
Even without an additional sag_amplitude loss, fine-tuning on the restricted non-firing regime of the dataset improves both trace prediction and sag_amplitude performance, since in this setting every sample is subthreshold and the feature can be computed consistently.
Including the feature-specific loss provides a further improvement of approximately 10% on sag_amplitude, while maintaining overall signal fidelity. Note that both settings converged after 300 epochs, ensuring that the comparison is fair.
These results demonstrate that prioritizing a feature through the loss function can yield targeted improvements without sacrificing global accuracy.
Appendix E Comparison of Sample Generation Time
The trained NOBLE generates predictions significantly faster than the reference numerical solver from [44].
We record the time necessary to generate 10,000 predictions on a workstation equipped with a single NVIDIA RTX 4090 GPU (24GB VRAM), an AMD Ryzen 9 7900X CPU, and 64GB of system RAM. The numerical solver only generates a single prediction at a time and takes roughly 36,200 seconds to generate 10,000 predictions. When doing one inference at a time with NOBLE (i.e. batch size of 1), we generate 10,000 predictions in 157 seconds, i.e. a speedup of approximately 230× compared to the solver.
In addition, NOBLE can easily be accelerated on a single GPU by generating multiple predictions at the same time. In particular, with a batch size of 1000, NOBLE generates 10,000 predictions in 8.59 seconds, i.e. a speedup of approximately 4200× compared to the solver.
Appendix F Scope and Future Directions
NOBLE successfully captures the dynamics across HoF models but is currently limited to single-neuron settings. A natural next step is to extend NOBLE to multi-neuron configurations with time-varying stimuli, enabling applications such as neuron classification and predicting multi-neuron dynamics. The embedding space used in our experiments is low-dimensional, constructed from two interpretable features derived from a biological neuron model’s F-I curve. Extending this to a learnable, higher-dimensional continuous embedding space represents a promising direction for future work. Beyond this, NOBLE could also integrate additional modalities such as gene expression or morphology to build a unified latent space linking molecular, electrophysiological, and structural characteristics, as discussed in the main text.
Although NOBLE demonstrates strong performance in modeling nonlinear neuronal dynamics, our study makes a few deliberate scope choices. These do not represent inherent limitations of the framework but rather natural starting points, each of which can be extended with minimal or no modifications.
  • Input currents: We restricted our attention to square-pulse DC step currents as these are widely used in electrophysiological experiments and represent a common protocol for characterizing neuron behavior. However, the NOBLE framework is not specific to these types of input currents: time-varying inputs can be incorporated directly by including examples during training. In such cases, the amplitude embedding can be removed or adapted (e.g., embedding a function of the amplitude, such as a moving-average modulation or the maximum amplitude).
  • Choice of operator learning architecture: We used the FNO primarily for its computational efficiency and strong generalizability. Moreover, since FNOs use the FFT and thus require inputs and outputs on equidistant grids, they align naturally with our data, where both simulations and experimental recordings are sampled at constant timesteps. For non-uniformly spaced data, alternative neural operators such as geometry-informed neural operators (GINOs) [63, 58] could be used within the same framework.
  • Training cost vs. efficiency: Training NOBLE on 75,600 samples for 300 epochs took approximately four days on a 64GB NVIDIA Tesla P100 GPU, which is small compared to the $\sim$600,000 CPU hours required to generate the HoF models via evolutionary optimization. However, once trained, NOBLE enables fast inference and the instantaneous generation of infinitely many bio-realistic voltage traces through latent space interpolation, capabilities not possible with the original PDE models.
  • Neuron populations considered: We focused on inhibitory neurons (PVALB and VIP), which show strong heterogeneity in morphology, gene expression, and electrophysiology, making them a stringent test for generalization. However, NOBLE is not restricted to inhibitory neurons and can naturally extend to excitatory types. It can also be applied to larger populations or multiple neurons within a family by expanding the latent embedding space with additional electrophysiological features that capture intracellular variability.